Calculate the mass of boron required to make a silicon crystal with 10^16
/cm^3
doping density, if the initial melt load of silicon is 50 kg. The density of silicon in
the melt is 2.5 g/cm^3
and boron has an atomic weight of 10.8 u. Assume that the
equilibrium segregation coefficient is constant throughout the growth process.
1
Expert's answer
2020-02-17T09:17:19-0500

The density of silicon in the melt is 2.5 g/cm3

Volume of the melt = mass/density = 50000/2.5 = 20000 cm3


Number of particles of boron in 20000 cm3 ="10^{16}*20000=2*10^{20}"

1 mole= "6.02*10^{23}" particles

x moles ="2*10^{20}" particles

x="\dfrac{2}{6.02}*10^{-3}" moles

Mass of boron in x moles= "\dfrac{1}{3}*10^{-3}*10.8=" "3.6*10^{-3}grams"